Two objects are connected by a very light and flex
Answer #1

a) FBD of the masses

m1 m2 m1g Free Body Diagrams of m, and m mi

b) m2g - T = m2a

T - m1g = m1a

adding both equations we have,

m2g-m1g = a(m1+m2)

a = g(m2-m1) / (m1 + m2)

a = 9.8*(3.27 - 1.25) / (3.27+1.25)

a = 4.38 m/s2

c) tension in the string

T = m1(g+a)

T = 1.25(9.8+4.38)

T = 17.72 N